How many grams of magnesium sulfate are needed to make 250.0 mL of a 0.250 M solution? my response was a wrong answer 7.5 grams

0.250M *(250mL* 1L/1000mL)= 0.0625

0.0625* 120.4 g/mol (MgSO4)= 7.525

7.5 grams
Answer #1

Molarity =Number of moles of solute/volume of solution in moles

Therefore, 0.250 M=number of moles of MgSO4/ 0.250 L

Therefore, Number of moles of of MgSO4=0.250 M*0.250 L= 0.0625 Moles of MgSO4

Now, Number of moles of MgSO4 =mass of MgSO4 /molar mass of MgSO4

Therefore, 0.0625 moles= mass of MgSO4 /120.366 (g/mol)

Therefore, mass of MgSO4 =0.0625*120.366 g=7.5229 grams

Thank You!